a car of mass M is moving with speed v. the brake of mass m and specific heat capacity c, is used to stop the car. if half of the K.E is absorbed by the brake thenwhat is the increase in the temperature of the brake


1
Expert's answer
2021-09-04T07:35:10-0400

Solution.

"W_k=\dfrac{Mv^2}{2};"

"Q=cm\Delta T \implies \Delta T=\dfrac{Q}{cm}" ;


"Q=\dfrac{W_k}{2};"

"\Delta T=\dfrac{Mv^2}{4cm};"

Answer: "\Delta T=\dfrac{Mv^2}{4cm}."